The DJ Company issued a 12-year semi-annual pay bond 2 years ago. The face value of the bond is $1,100 and the coupon rate is 4%. The current market rate (annual APR) on comparable bonds is 5%. Calculate the bond's current price.
Answer #1

n = 10 * 2 = 20 semi-annual periods, M = $1100, i = 5%/2 = 2.5% (semi-annually), C = 4% * $1100/2 = $22 (semi-annually)

P = 22* - (1+0.025)20 0.025 - + 1100 (1+0.025 20

P = $342.96 + $638.95

P = $981.91